if f(x)=2x-7 and g(x)=x^2-4x-5, find [gof](5)
Answer by stanbon(75887)

if f(x)=2x-7 and g(x)=x^2-4x-5, find [gof](5)
--------------
[gof](5) = [g[f(5)] = g[2*5-7]
= g[3] = 3^2-4*3-5 = -8
===================
Cheers,
Stan H.